The temperature of an aluminum disk is increased by 220 ∘C. By what percentage does its volume increase?
Answer #1

Solution)

Given

ΔT = 220 C

We know, β of alluminium = 69*10^-6

Now,

Change in volume ΔV/V = β * ΔT = 69*10^-6 * 220

ΔV/V = 0.01518

% of volume increase = ΔV/V * 100 = 1.518 % (Ans)

==========

Good luck!:)